OpenStudy (mendicant_bias):

I'm reviewing Cauchy-Euler ODE's and am going to be trying to solve the following problem, don't remember much about these. Posted below momentarily.

OpenStudy (mendicant_bias):

\[x^{2}y''-2y=0\]

OpenStudy (mendicant_bias):

If I remember correctly, the first thing we're supposed to do when solving this is solve the characteristic equation.

OpenStudy (mendicant_bias):

Is the characteristic equation supposed to be \[m^2 - 2m ?\]

OpenStudy (mendicant_bias):

@dan815

OpenStudy (mendicant_bias):

(Or, is this how I'm supposed to be approaching the problem?)

OpenStudy (dan815):

u get the characteristic by assuming a solution is in form y=x^n

OpenStudy (dan815):

\[y'=n*^{n-1}\\ y''=n(n-1)*x^{n-2}\\ subbing\\ x^2y^{''}-2y=0\\ x^2*n(n-1)*x^{n-2}-2*x^n=0\\ n(n-1)*x^n-2*x^n=0\\ x^n(n(n-1)-2)=0\\ x^n=0,n(n-1)-2=0\\ n^2-n-2=0\\ (n+1)(n-2)=0; n=-1,2\]

OpenStudy (dan815):

\[therfore~~ y=x^{-1}= 1/x , y=x^2\]

OpenStudy (dan815):

Y_gen=c1/x+c2x^2

OpenStudy (mendicant_bias):

Alright, cool, totally forgot how to do this, lol. Looking at this, one second.

OpenStudy (mendicant_bias):

Alright, yeah, this makes sense; I don't entirely understand the math behind x^n equaling zero and then it disappearing, but I get how to do this (?). Now I just need to remember how to do the rest of this. Now I find a particular solution?